OpenStudy (anonymous):

sketch the solid obtained by rotating the region underneath the graph of the function over the given interval about the y-axis, and find its volume. f(x) = sqrt(x), [0,4]

OpenStudy (dumbcow):

easiest to use shell method here radius is x height is f(x) \[V = 2\pi \int\limits_{0}^{4}x \sqrt{x} dx\]